# Learning control variables and instruments for causal analysis in observational data

### v1

@@ 4 Selection Method
We note that the first step is required to select candidate instruments which, if ultimately selected as instrument $Z$, are statistically significant in terms of satisfying the first stage condition $E[D|X,Z]\neq E[D|X]$. Let us assume that $Q$ is low dimensional in the sense that the sample size $n$ is larger than the number of variables in $Q$, denoted by $p$, such that $n>p$. In this case, step one might be implemented based on a first stage regression of $D$ on $Q$ and selecting all regressors with statistically significant associations after controlling for multiple hypothesis testing issues into the set of candidate instruments $\hat{S}$.
As statistical criterion for selecting candidate instruments, we suggest first-stage hard thresholding, as performed in the IV selection literature for instance by Guo et al. (2018) and Masten and Poirier (2021). To formalize the discussion, let $F_{r}$ be the first stage F-statistic in the estimation of $E[D|X,Z]$ when defining $Z=Q_{r}$ and $X=Q_{[r]}$, with $Q_{[r]}=Q\setminus Q_{r}$ (and $Q=X\cup Z$). We obtain the estimated set of strong candidate instruments, denoted by $\hat{S}$, as those whose F-statistic exceeds a certain critical value $C_{n}$:
$\hat{\mathcal{S}}=\{r:F_{r}\geq C_{n}\}$ (4.1)
@@
An alternative strategy is to iteratively consider each variable in $Q$ as instrument $Z$ and any remaining variables as covariates $X$ when estimating the first stage association $E[D|X,Z]$ based on methods for treatment or policy evaluation like doubly robust estimation, see e.g. Robins et al. (1994), Robins et al. (1995) and Chernozhukov et al. (2018). Such doubly robust approaches are also applicable in high-dimensional contexts with $n<<p$ when combining them with machine learning to control for important control variables among high-dimensional covariates $X$. For instance, the double machine learning algorithm suggested by Soleymani et al. (2022), selects direct causal parents in a data-driven manner by iteratively changing the definition of variables whose causal effect is of interest and the control variables. It may be applied to select the set of candidate instruments $S$ which are those variables in $Q$ strongly associated with $D$ conditional on the remaining elements in $Q$. Second, we proceed to test the null hypothesis in equation (3.2) iteratively over all candidate instruments that pass the first-stage threshold, $\hat{\mathcal{S}}$.
Our aim is to find a partition of variables for which the conditional independence of the respective candidate instrument holds. To discuss this more formally, we introduce the partition
@@
such that variable $j$ in set $Q$ is chosen to be the instrument, while the remaining variables in set $Q$ are used as control variables. Furthermore, let $\mathcal{V}$ denote the set of candidate instruments which are conditionally mean independent of the outcome, satisfying
@@
Furthermore, we denote the set of partitions for which the instrument has a first stage effect on the treatment and the conditional independence of the instrument holds by $\mathcal{P}^{*}$:
@@
$\hat{\mathcal{P}}_{pass}=\left\{\mathcal{P}_{j}:j\in\left(\hat{\mathcal{S}}\cap\hat{\mathcal{V}}\right)\right\},$ (4.3)
where $\hat{\mathcal{V}}=\{j:|\sqrt{n}\hat{\sigma}^{-1}_{j}\hat{\theta}_{j}|<c_{\alpha}\}$ is the set of instruments for which conditional independence is not rejected based on the test defined in Theorem 1 and Corollary 1 with significance level $\alpha$ and the corresponding critical value $c_{\alpha}$.
If $|\hat{\mathcal{P}}_{pass}|>1$ such that there is more than one single candidate instrument for which the first stage is sufficiently strong while conditional independence is not rejected, we can proceed in two ways. For a set of partitions $\mathcal{P}_{pass}=\{\mathcal{P}_{1},...,\mathcal{P}_{|\mathcal{P}_{pass}|},\}$ where each partition is associated with a specific IV, we define the set of IVs associated with that set of partitions as
=$\mathcal{I}=\left\{j:\mathcal{P}_{j}\in\mathcal{P}_{pass}\right\},$ (4.4)
@@
$\mathcal{P}_{all}=\{Z=Q_{\mathcal{I}},\ \ X=Q\setminus Q_{\mathcal{I}}\}.$ (4.5)
Alternatively, we can select the final partition as the one which maximizes the p-value when testing conditional independence:
$\hat{\mathcal{P}}_{pmax}=\underset{\mathcal{P}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\hat{\mathcal{P}}_{pass}),$ (4.6)
where $p(\mathcal{P})$ is the $p$-value obtained in the conditional independence test of the instrument and the outcome. Even more crucial than the question how to choose the final partition if $|\hat{\mathcal{P}}_{pass}|>1$ is the question whether $\hat{\mathcal{P}}_{pass}=\emptyset$ or not.
The main contribution of this paper is the development of a new statistical procedure which tests the identification of a causal effect in a data-driven way. To this end, we consider the following hypotheses:
$H_{0}:\text{no identification}\quad v.s\quad H_{1}:\text{identification}\quad(Y(d){\perp\!\!\!\perp}D|X\quad\forall d\in\mathcal{D}).$ (4.7)
We conclude that $H_{1}$ is true, i.e, our causal effect of interest is identified if $\hat{\mathcal{P}}_{pass}\neq\emptyset$. The following Theorem 2 helps us to understand the type 1 error of our proposed test.
@@
If $\hat{\mathcal{P}}_{pass}=\emptyset$, our statement holds trivially. Hence, consider a variable $j\in\hat{\mathcal{P}}_{pass}$, i.e, $Z_{j}\in\hat{\mathcal{S}}$ and $Z_{j}\in\hat{\mathcal{V}}$. Therefore, it holds
@@
4:     Regress: D $\leftarrow$ Q
5:     $\hat{F}_{j}\leftarrow\hat{t}_{j}^{2}$
6:     if $\hat{F}_{j}>C_{0.1/log(n)}$ then $j\in\hat{\mathcal{S}}$
7:     Select instruments passing conditional independence test $\hat{\mathcal{V}}$:
8:     for $j\in\hat{\mathcal{S}}$ do
9:         $\hat{\mu}(D,X,Z)\leftarrow$ and $\hat{m}(D,X)\leftarrow$ and $\hat{p}(D,X)\leftarrow$ estimates from machine learning
10:         $\hat{\theta}\leftarrow$ constructed via score in equation (3.4) (binary case) or (3.5) (continuous case)
11:         $\hat{t}_{ind,j}\leftarrow$ test $H_{0}:\hat{\theta}=0$
12:         if $\hat{t}_{ind,j}<t_{crit}$ then $j\in\hat{\mathcal{V}}$
13:     Select final partition $\hat{\mathcal{P}}$:
14:     $\hat{\mathcal{P}}_{pass}\leftarrow\{j\in\hat{\mathcal{S}},j\in\hat{\mathcal{V}}\}$
15:     if $\hat{\mathcal{P}}_{pass}=\emptyset$ then end and report that identification is “rejected”
16:     else if $|\hat{\mathcal{P}}_{pass}|=1$ then $\hat{\mathcal{P}}\leftarrow\hat{\mathcal{P}}_{pass}$
17:     else if $|\hat{\mathcal{P}}_{pass}|>1$ then
18:         $p(\hat{\mathcal{P}}_{pass})\leftarrow\text{p-value}(\hat{t}_{ind,j})$ for $j\in\hat{\mathcal{V}}$
19:         $\hat{\mathcal{P}}=\underset{\mathcal{P}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\mathcal{P})$
20:     In case identification is confirmed ($\hat{\mathcal{P}}_{pass}\neq\emptyset$), estimate the treatment effect:
21:     Regress: $Y\leftarrow D$ and $X=Q_{[j]}\in\hat{\mathcal{P}}$
@@ 5 Simulation study
with $X,Z,W,U,V$ being independent of each other. Outcome $Y$ is a linear function of $D$ (whose treatment effect is one), covariates $X$ (for $\beta\neq 0$), the unobservables $W$ and $U$, and the supposed instrument $Z$ if the coefficient $\gamma\neq 0$. The binary treatment $D$ is a function of $X$ and the unobservable $V$, as well as $W$ if coefficient $\delta\neq 0$. While the supposed instrument $Z$ is binary, the unobserved terms $U,V,W$ are random, standard normally distributed variables that are independent of each other, of $Z$, and of $X$. Covariates $X$ are created as follows. We first draw a matrix of multivariate normal variables $\tilde{X}\sim N(\mathbf{0},\bm{\Sigma})$ where $\bm{\Sigma}$ is a covariance matrix with $Cov(X_{j},X_{k})=0.5^{|j-k|}$. We then compute probabilities $\pi_{i}=\frac{1}{1+exp(-\tilde{X}_{i})}$ and draw $X_{i}\sim Bernoulli(\pi_{i})$. $\beta$ gauges the effects of the covariates on $Y$ and $D$, respectively, and thus, the magnitude of confounding due to observables. The $j$th element in the coefficient vector $\beta$ is set to $0.8/j$ for $j=1,...,4$, implying a linear decay of covariate importance in terms of confounding for the first 4 covariates. We use the final element, $j=20$ as the instrument $Z$. This results in 20 total covariates, with the first 4 as confounders, and the last as the instrument $Z$.
@@
We investigate the performance of our testing approach in $100$ simulations with sample sizes of $n=1000$, $4000$, and $16000$, setting the number of covariates to $20$. Testing is based on the cross-fitted estimator $\hat{\theta}$ of equation (3.6) with two folds ($K=2$) when using the doubly robust score function in equation (3.4). For nuisance parameter estimation, lasso regression with default parameters as implemented in the glmnet package by Friedman et al. (2010) for the statistical software R is applied. We consider several statistics from our simulations: the estimated violation, $\hat{\theta}$, when using variable $Z$ as the instrument (est), the standard deviation of these estimates (std), and the average of the standard errors across all samples (mean se). These statistics are useful for judging the performance of the test conditional on selecting $Z$ as the instrument, which is the correct choice if $\gamma=0$, while there is no valid instrument if $\gamma\neq 0$.
@@
However, since the choice of the instrument is not fixed a priori but is part of the estimation process, we also consider the following two performance measures. The first one is the empirical detection rate of $Z$ (det.Z), which indicates the frequency with which variable $Z$ is selected as the best candidate instrument and has a p-value greater than 10% when testing conditional independence based on the estimate $\hat{\theta}$. In other words, this corresponds to the proportion of simulations in which our method simultaneously selects $Z$ as the instrument and keeps the null hypothesis. In scenarios where the null hypothesis holds ($\delta=\gamma=0$), this detection rate should approach one as the sample size increases. The second measure is an equivalent detection rate for the confounders $X_{1}$ to $X_{4}$ (det.X). It indicates the frequency with which one of the confounders is selected as the best candidate instrument and yields a p-value greater than 10% when testing conditional independence using this confounder as the instrument. As the sample size increases, this detection rate should always approach zero, because confounders are never valid instruments.
Table 1: Simulations: Binary Instrument
N est std mean se det.Z det.X
@@
1000 -0.353 0.185 0.146 6% 69%
4000 -0.457 0.111 0.091 0% 72%
16000 -0.578 0.072 0.053 0% 2%
@@
Notes: columns ‘est’, ‘std’, and ‘mean se’ provide the average estimate of $\theta$ (the violation of conditional independence) conditional on using $Z$ as instrument, its standard deviation, and the average of the standard errors across all samples, respectively. ‘det.Z’ gives the empirical detection rate of selecting $Z$ as instrument and obtaining a p-value $>0.10$ (or 10%) when estimating $\theta$. This corresponds to the frequency with which $Z$ is selected as the best candidate instrument and also entails a p-value of more than 10% when testing conditional independence. ‘det.X’ is an equivalent measure for confounders $X_{1}$ to $X_{4}$. It corresponds to the frequency with which one of the confounders is selected as the best candidate instrument and also entails a p-value of more than 10% when testing conditional independence.
@@
Table 1 reports the simulation results for a binary $Z$ and binary $X_{j}$. $Z$ is correlated with $X_{j}$ according to the covariance matrix $\bm{\Sigma}$. The top rows focus on the case where $\delta=\gamma=0$, meaning that both Assumptions 4 and 5 are satisfied. As the sample size increases, the test is more likely to simultaneously select the true instrument $Z$ and keep the null hypothesis that the assumptions are satisfied. Specifically, the detection rate of $Z$ (det.Z) increases from 31% with the smaller sample size of 1000 observations to 61% with 16000 observations. This improvement is mirrored by a reduction in the rate at which a confounder variable ($X_{1}...X_{4}$) is selected as the instrument and passes the test (det.X). Under the largest sample size of 16000 observations, it never occurs that any confounder is selected as the best instrument and passes the conditional independence test. Additionally, we observe that, conditional on having selected the true instrument $Z$, the estimated violation $\hat{\theta}$ (est) is close to zero for any sample size.

### v2

@@ 4 Selection Method
We note that the first step is required to select candidate instruments which, if ultimately selected as instrument $Z$, are statistically significant in terms of satisfying the first stage condition $E[D|X,Z]\neq E[D|X]$. Let us assume that $Q$ is low-dimensional in the sense that the sample size $n$ is larger than the number of variables in $Q$, denoted by $p$, $n>p$. In this case, step one might be implemented based on a first stage regression of $D$ on $Q$ and selecting all regressors with statistically significant associations after controlling for multiple hypothesis testing issues into the set of candidate instruments $\hat{S}$. In high-dimensional settings where $p>n$, lasso based approaches can be applied to select strong IV candidates. Details are given in the next subsection.
4.1 Strong IV Selection
To select strong instrumental variables (IVs) in the high-dimensional setting, we use first-stage hard-thresholding (FSHT) proposed by Guo et al. (2018). In this approach, instruments are considered irrelevant if their t-statistic does not exceed a predefined threshold. [^8] Specifically, an instrument is treated as weak if $|t_{\gamma_{j}}|=\left|\frac{\hat{\gamma}_{j}}{\sqrt{\widehat{Var}(\hat{\gamma}_{j})}}\right|<\sqrt{2.01\log(\max(p,n))}.$ (4.1) Here, $\widehat{Var}(\hat{\gamma}_{j})$ denotes the variance estimator of the first-stage coefficient estimator $\hat{\gamma}_{j}$. The right-hand side accounts for multiple testing, where the factor $\log(\max(p,n))$ originates from the tail bound of the normal distribution. In the low-dimensional case, this term simplifies to $\sqrt{2.01\log(n)}$. This procedure is shown to consistently select relevant instruments, as demonstrated in Lemma 1 in Guo et al. (2018). Windmeijer et al. (2021) treat the factor 2.01 as a tuning parameter and show in simulations that this choice of the tuning parameter performs well in practice. Meanwhile, Guo et al. (2018) argue that the consistent selection property holds as long as the argument inside the logarithm grows with $n$. We iteratively consider each variable in $Q$ as instrument $Z$ and any remaining variables as covariates. That is, when considering the $j$th variable in $Q$ as instrument and denoting it by $Q_{j}$, we have that $Z=Q_{j}$ and $X=Q_{[j]}$, where $Q_{[j]}=Q\setminus Q_{j}$ and $Q=X\cup Z$ when estimating the first-stage association $E[D|X,Z]$. Leveraging the double machine learning (DML) literature, we consider the following partially linear regression (PLR) model specification to estimate the effect of a variable $Z$ on treatment $D$:
$D$ $=\gamma_{j}^{T}Z+g(X)+\varepsilon,$ (4.2)
$Z$ $=h(X)+v.$ (4.3)
Here, $g(X)$ and $h(X)$ are general functions of $X$, and the first-stage effect $\gamma$ of $Z$ is indexed by $j$ to emphasize that it may vary with each variable considered as an instrument. Note that although the partial linear model assumes additive separability and a linear effect of $Z$ on $D$, estimation based on this model may also have power in fully nonparametric settings, where the effect of $Z$ on $D$ may be nonlinear and may interact with $X$. In such cases, it follows from Stein’s Lemma (Stein, 1981) that $\gamma_{j}$ corresponds to a weighted average of the marginal effects of $Z$ on $D$ given $X$, with all weights being nonzero. [^9] In special cases (in particular $Z$ being Gaussian conditional $X$), $\gamma_{j}$ is the (unweighted) average of marginal effects.
DML-based estimation of $\gamma_{j}$ is asymptotically normal under certain regularity conditions (in particular, when estimators of the models of $D$ and $Z$ attain a convergence rate of $o(n^{-1/4})$), see Chernozhukov et al. (2018): $\sqrt{n}\hat{\sigma}_{\gamma,j}^{-1}(\hat{\gamma}_{j}-\gamma_{0,j})\overset{d}{\rightarrow}N(0,1)$, where $\hat{\sigma}_{\gamma,j}^{-1}$ is the standard error of $\hat{\gamma}_{j}$. As only one IV is considered in turn, the first-stage F-statistic is equal to the square of the t-statistic, $t_{j}^{2}=F_{j}$, which is asymptotically $\chi_{1}^{2}$-distributed with one degree of freedom, $F_{j}\overset{d}{\rightarrow}\chi_{1}^{2}$ (see Proposition 3 of Masten and Poirier (2021)). Under the alternative hypothesis, $\gamma_{j}\neq 0$, we have $\frac{F_{j}}{n}\overset{P}{\rightarrow}\kappa_{j}$, where $\kappa_{j}>0$ is a constant. The critical values $C_{n}$ for the F-statistic need to satisfy $C_{n}\to\infty$ and $C_{n}=o(n)$ as $n\to\infty$, for the first-stage hard-thresholding procedure to successfully identify strong candidate instruments with probability approaching 1, i.e.,
$\lim\limits_{n\rightarrow\infty}P(\hat{S}=\mathcal{S})=1,$
@@ 4.1 Strong IV Selection
4.2 Valid IV Selection and Identification Test
In the second step, we test the null hypothesis in equation (3.2) iteratively over all candidate instruments that pass the first-stage threshold, $\hat{\mathcal{S}}$. Our aim is to find a partition of variables for which the conditional independence of the respective candidate instrument holds. To discuss this more formally, we introduce the partition
@@ 4.2 Valid IV Selection and Identification Test
such that variable $j$ in set $Q$ is chosen to be the instrument, while the remaining variables in set $Q$ are used as control variables. Furthermore, let $\mathcal{V}$ denote the set of candidate instruments which are conditionally mean independent of the outcome, thus satisfying condition (2.2) and the null hypothesis in equation (3.2):
@@
Moreover, we denote the set of partitions for which the instrument has a first stage effect on the treatment and the conditional independence of the instrument holds by $\mathcal{P}^{*}$:
@@
[…]
where $\hat{\mathcal{V}}=\{j:|\sqrt{n}\hat{\sigma}^{-1}_{j}\hat{\theta}_{j}|<c_{\alpha}\}$ is the set of instruments for which conditional independence is not rejected based on the test defined in Theorem 1 and Corollary 1 with significance level $\alpha$ and the corresponding critical value $c_{\alpha}$. If $|\hat{\mathcal{P}}_{pass}|>1$ such that there is more than one single candidate instrument for which the first stage is sufficiently strong while conditional independence is not rejected, we can proceed in two ways. For a set of partitions $\hat{\mathcal{P}}_{pass}=\{\mathcal{P}_{1},...,\mathcal{P}_{|\hat{\mathcal{P}}_{pass}|},\}$ where each partition is associated with a specific IV, we define the set of IVs associated with that set of partitions as
$\hat{\mathcal{I}}=\left\{j:\mathcal{P}_{j}\in\hat{\mathcal{P}}_{pass}\right\},$ (4.6)
@@
$\hat{\mathcal{P}}_{all}=\{Z=Q_{\hat{\mathcal{I}}},\ \ X=Q\setminus Q_{\hat{\mathcal{I}}}\}$. (4.7)
In this paper, we select the final partition as the one which maximizes the p-value when testing conditional independence:
$\hat{\mathcal{P}}_{pmax}=\underset{\mathcal{P}_{j}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\mathcal{P}_{j}),$ (4.8)
where $\texttt{p}(\mathcal{P}_{j})$ is the p-value obtained in the conditional independence test of the instrument and the outcome. Even more crucial than the question how to choose the final partition if $|\hat{\mathcal{P}}_{pass}|>1$ is the question whether $\hat{\mathcal{P}}_{pass}=\emptyset$ or not. The main contribution of this paper is the development of a new statistical procedure which tests the identification of a causal effect in a data-driven way. To this end, we consider the following hypotheses:
$H_{0}:\text{no identification}\quad vs.\quad H_{1}:\text{identification (conditional mean independence}).$ (4.9)
We conclude that $H_{1}$ is true (the ATE is identified, see the null hypothesis in (3.1)) if $\hat{\mathcal{P}}_{pass}\neq\emptyset$. The following Theorem 2 helps us to understand the type 1 error of our proposed test.
@@
If $\hat{\mathcal{P}}_{pass}=\emptyset$, our statement holds trivially. Hence, consider a variable $j:\mathcal{P}_{j}\in\hat{\mathcal{P}}_{pass}$, i.e, $Z_{j}\in\hat{\mathcal{S}}$ and $Z_{j}\in\hat{\mathcal{V}}$. Therefore, it holds
@@
4:  for $j\in Q$ do
5:   Predict $D$ by $Q_{j}$ conditional on $Q\setminus Q_{j}$ using machine learning or regression
6:   $\hat{F}_{j}\leftarrow\hat{t}_{j}^{2}$
7:   if $\hat{F}_{j}>C_{\tilde{\alpha}}$ with $\tilde{\alpha}=0.1/log(n)$ $j\in\hat{\mathcal{S}}$ then
8:  Select instruments passing conditional independence test $\hat{\mathcal{V}}$:
9:  for $j\in\hat{\mathcal{S}}$ do
10:   $\hat{\mu}(D,X,Z)\leftarrow$ and $\hat{p}(D,X)\leftarrow$ estimates from machine learning
11:   $\hat{\theta}_{j}\leftarrow$ constructed via score in equation (3.4) (binary case) or (3.5) (continuous case)
12:   $\hat{\sigma}_{j}\leftarrow$ constructed as described in Theorem 1
13:   $\hat{t}_{ind,j}=\frac{\hat{\theta}_{j}}{\hat{\sigma}_{j}}\leftarrow$ test $H_{0}:\theta=0$
14:   if $\hat{t}_{ind,j}<c_{\alpha}$ then $j\in\hat{\mathcal{V}}$
15:  Select final partition $\hat{\mathcal{P}}$:
[…]
17:  if $\hat{\mathcal{P}}_{pass}=\emptyset$ then end and report that $H_{0}$ is not rejected (no identification)
18:  else if $|\hat{\mathcal{P}}_{pass}|=1$ then $\hat{\mathcal{P}}\leftarrow\hat{\mathcal{P}}_{pass}$
19:  else if $|\hat{\mathcal{P}}_{pass}|>1$ then
20:   $\texttt{p}(\mathcal{P}_{j})\leftarrow\texttt{p}-value(\hat{t}_{ind,j})$ for $j\in\hat{\mathcal{V}}$
21:   $\hat{\mathcal{P}}_{pmax}=\underset{\mathcal{P}_{j}\in\hat{\mathcal{P}}_{pass}}{argmax}\ \ p(\mathcal{P}_{j})$
22:  In case identification is confirmed ($\hat{\mathcal{P}}_{pass}\neq\emptyset$), estimate the treatment effect:
23:  Regress: $Y\leftarrow D$ and $X=Q_{[j]}\in\hat{\mathcal{P}}_{pmax}$
@@ 5 Simulation study
with $X,Z,W,U,V$ being independent of each other. Outcome $Y$ is a linear function of $D$ (whose treatment effect is one), covariates $X$ (for $\beta\neq 0$), the unobservables $W$ and $U$, and the supposed instrument $Z$ if the coefficient $\gamma\neq 0$. The binary treatment $D$ is a function of $X$ and the unobservable $V$, as well as $W$ if coefficient $\delta\neq 0$. While the supposed instrument $Z$ is binary, the unobserved terms $U,V,W$ are random, standard normally distributed variables that are independent of each other, of $Z$, and of $X$. The joint set of covariates and instruments $Q$ is created as follows. We first draw a matrix of multivariate normal variables $\tilde{Q}\sim N(\mathbf{0},\bm{\Sigma})$ where $\bm{\Sigma}$ is a covariance matrix with $Cov(\tilde{Q}_{j},\tilde{Q}_{k})=0.5^{|j-k|}$. We then compute probabilities $\pi_{i}=\frac{1}{1+exp(-2\cdot\tilde{Q}_{i})}$ and draw $Q_{i}\sim Bernoulli(\pi_{i})$. There are in total 10 variables in the set of observed variables $Q=(X,Z)$, with the last one being the (single) instrument $Z$. Concerning the first nine elements in $Q$, which are considered as covariates $X$, $\beta$ gauges their effects on $Y$ and $D$, respectively, and thus, the magnitude of confounding due to observables. The $j$th element in the coefficient vector $\beta$ is set to $0.8/j$ for j=1,,4j=1,...,4, implying a linear decay of covariate importance in terms of confounding for the first 4 covariates. Furthermore, for elements j=5,,9j=5,...,9, the coefficients are equal to zero, implying that covariates 5 to 9 are random noise variables that are neither confounders, nor instruments.
In a second simulation design, we also consider continuous variables $Q$, based on the following data generating process:
$Y$ $=$ $D+X^{\prime}\beta+\gamma Z+W+U,$
$D$ $=$ $I\{X^{\prime}\beta+Z+\delta W+V>0\},$
$X,Z$ $\sim$ $Uniform(-0.5,0.5),$
$W$ $\sim$ $N(0,1),U\sim N(0,1),V\sim N(0,1).$
The set of observed covariates $X$ and the instrument $Z$ are now drawn from the multivariate uniform distribution, $X,Z\sim Uniform(-0.5,0.5)$. All other properties of this data generating process are identical to the previous scenario with binary $Z$ and $X$ variables. Specifically, the outcome $Y$ is characterized by the same linear function of $D,X,W$, and $U$ (and the supposed instrument $Z$ in the case where $\gamma\neq 0$) as before, and we again consider 10 variables in the set $Q=(X,Z)$, with the last one being the (single) instrument $Z$, elements j=1,,4j=1,...,4 being observed confounders, and elements j=5,,9j=5,...,9 being random noise variables that are neither confounders nor instruments.
@@
We investigate the performance of our testing approach in $200$ simulations with sample sizes of $n=1000$, 40004000, and $16000$. For estimating the (conditional) first stage effect of elements in $Q$ on $D$ to select sufficiently strong candidate instruments, we use Double Machine Learning (DML) for partially linear models (Chernozhukov et al., 2018). DML can be applied to both discrete and continuous elements in $Q$ and is implemented in the DoubleML package for statistical software R. The so-called nuisance parameters in the first stage are estimated using lasso regression with cross-fitting, setting the number of folds to five (K = 5). Specifically, the classification lasso is used for discrete elements in $Q$, modelling probabilities for treatment assignment while lasso regression is used for continuous variables. For testing the conditional independence of any selected candidate instrument in the second stage, we consider a cross-fitted estimator $\hat{\theta}$ of equation (3.6) with five folds ($K=5$), which is based on the doubly robust score functions in equations (3.4) and (3.5) for binary and continuous candidate instruments, respectively. In the case of a continuous candidate instrument, its support is partitioned based on the quartiles of its distribution, such that $L=4$. For nuisance parameter estimation in the second stage, lasso regression with default parameters as implemented in the glmnet package by Friedman et al. (2010) for the statistical software R is applied. We consider several statistics from our simulations: the estimated violation, $\hat{\theta}$, when using variable $Z$ as the instrument, the standard deviation of the estimated violation (std), and the average of the standard error of the estimated violation across all simulation samples (mean se). These statistics are useful for judging the performance of the test conditional on selecting $Z$ as the instrument, which is the correct choice if $\gamma=0$, while there is no valid instrument if $\gamma\neq 0$ and no identification when controlling for covariates if $\delta\neq 0$.
@@
Since the choice of the instrument is not fixed a priori but is part of the estimation process, we also consider three further statistics. The first one is the empirical selection rate of $Z$ (sel.Z), which indicates the frequency with which variable $Z$ is selected as the best candidate instrument and also has a p-value greater than 30% when testing conditional independence based on the estimate $\hat{\theta}$. [^10] We suggest the threshold of 30% based on examination of the distribution of p-values for both true instruments and confounders, see Figures A1 and A2. This stricter threshold guards against the possible choice of a confounder as an instrument, although a more traditional threshold of 10% could also be used. In other words, this corresponds to the proportion of simulations in which our method simultaneously selects $Z$ as the instrument and at the same time keeps the null hypothesis of conditional independence. In scenarios where the null hypothesis holds ($\delta=\gamma=0$), this selection rate should approach one as the sample size increases. The second statistic is the analogous selection rate of the noise variables $X_{5}$ to $X_{9}$ (sel.noconf). It indicates the share of simulations in which one of the non-confounders (that are not instruments either) is selected as the best candidate instrument and at the same time yields a p-value greater than 30% when testing conditional independence when using this noise variable as the instrument. As the sample size increases, this selection rate should always approach zero, because asymptotically, noise variables do not have a first stage effect on the treatment conditional on other elements in $Q$. Therefore, noise variables should not be selected as candidate instrument for the conditional independence test. Finally, we report an equivalent selection rate for the confounders $X_{1}$ to $X_{4}$ (sel.conf). It corresponds to the share of simulations in which one of the confounders is selected as the best candidate instrument and at the same time yields a p-value greater than 30% when testing conditional independence when using this confounder as instrument. As the sample size increases, this selection rate should always approach zero, because confounders are never valid instruments.
Table 1: Simulations: Single Instrument
PANEL A: Binary Instrument
N sel.Z sel.noconf sel.conf $\hat{\theta}$ std mean se
@@
1000 30% 2% 57% 0.045 0.040 0.050
4000 3% 2% 68% 0.053 0.019 0.027
16000 0% 3% 8% 0.054 0.010 0.014
PANEL B: Continuous Instrument
N sel.Z sel.noconf sel.conf $\hat{\theta}$ std mean se
Assumptions 4 and 5 hold ($\delta=0$, $\gamma=0$)
1000 55% 2% 19% -0.007 0.029 0.031
4000 65% 4% 8% -0.001 0.015 0.016
16000 64% 1% 0% 0.000 0.007 0.009
Ass. 4 violated, Ass. 5 holds ($\delta=2$, $\gamma=0$)
1000 16% 6% 6% -0.018 0.031 0.017
4000 8% 2% 1% -0.015 0.018 0.007
16000 0% 2% 0% -0.030 0.014 0.003
Ass. 4 holds, Ass. 5 violated ($\delta=0$, $\gamma=0.5$)
1000 0% 4% 18% -0.490 0.192 0.049
4000 0% 6% 8% -0.552 0.103 0.026
16000 0% 1% 0% -0.554 0.055 0.013
@@
Notes: columns $\hat{\theta}$’, ‘std’, and ‘mean se’ provide the average estimate of $\theta$ (the violation of conditional independence) conditional on using $Z$ as instrument, its standard deviation, and the average of the standard errors across all samples, respectively. ‘sel.Z’ gives the share of simulations in which $Z$ is selected as best candidate instrument and simultaneously yields a p-value >0.30>0.30 (or 30%) when testing conditional independence based on the estimate of $\hat{\theta}$. ‘sel.noconf’ is an equivalent measure for covariates $X_{5}$ to $X_{9}$, corresponding to the share of selecting a non-confounder as best candidate instrument and having a p-value >> 30% when testing conditional independence. ‘sel.conf’ is an equivalent measure for covariates $X_{1}$ to $X_{4}$, corresponding to the share of selecting an observable confounder the best candidate instrument and having a p-value >> 30% when testing conditional independence.
Figure 2: Simulations: Distribution of Estimates
This figure depicts density plots of the estimated violation $\hat{\theta}$ for the true IV (covariate $X_{10}$), across the simulation settings. The top and bottom rows are the binary and continuous IV cases, respectively. The first plots on the left provide results for $\delta=\gamma=0$, i.e. Assumptions 4 and 5 are satisfied. The middle plots depict the scenario of Assumption 4 being violated, and the right-most plots Assumption 5 is violated. Orange is the smallest sample size (N = 1000), green the medium sample size (N=4000), and purple is the largest (N=16000).
@@
Table 1 reports the simulation results for binary variables in $Q$ in Panel A, and for continuous elements in $Q$ in Panel B, respectively. In both panels, the top rows provide the results for $\delta=\gamma=0$, implying that both Assumptions 4 and 5 are satisfied and conditional independence holds. As the sample size increases, the test is more likely to simultaneously select the true instrument $Z$ and keep the null hypothesis of conditional independence. Specifically, the selection rate of $Z$ (sel.Z) increases from 32% with the smaller sample size of 1000 observations to 68% with 16000 observations for the binary instrument, and from 55% with 1000 observations to 64% with 16000 observations for the continuous instrument. This improvement is mirrored by a reduction in the rate at which a confounder variable ($X_{1}...X_{4}$) is selected as the instrument and passes the test (sel.conf). Under the largest sample size, it only occurs that any confounder is selected as the best instrument and passes the conditional independence test 4% of the time for the binary and never for the continuous scenario. However, we observe that the noise variables ($X_{5}...X_{9}$) sometimes appear to step in, as their selection rate (sel.noconf) may transiently increase in response to the decreasing selection rate of confounders. Additionally, we observe that, conditional on having selected the true instrument $Z$, the estimated violation $\hat{\theta}$ is close to zero for any sample size and never statistically significant at conventional significance levels. Figure 2 depicts the distribution of the estimated $\hat{\theta}$ and visually confirms this finding in the first column. The top row of density plots show the binary $\hat{\theta}$ and the bottom row shows the continuous estimates. As sample size increases, the distribution moves closer to zero, as expected.